OpenStudy (anonymous):

Which numbers should be multiplied to obtain 1832 - 1392? 44 and 44 1936 and 1936 44 and 322 33350 and 33628

OpenStudy (faolan):

i dont get the question...

OpenStudy (faolan):

44 times 44 is 1936

OpenStudy (welshfella):

neither do I

OpenStudy (anonymous):

183^2 - 139^2 is what is says, but for some reason it keeps changing it

ganeshie8 (ganeshie8):

Hint : \(\large a^2-b^2 ~~=~~ (a-b)(a+b)\)

OpenStudy (anonymous):

Okay, so then (183-139)(183+139)

ganeshie8 (ganeshie8):

Yes

ganeshie8 (ganeshie8):

183-139 = ? 183+139 = ?

OpenStudy (anonymous):

44 322

OpenStudy (faolan):

Deez Nuts!!

OpenStudy (faolan):

jk 44 and 44

OpenStudy (faolan):

nope

OpenStudy (anonymous):

So its C then

OpenStudy (faolan):

srry subtracted both ur right

OpenStudy (anonymous):

Thanks, I completely forgot about that formula

ganeshie8 (ganeshie8):

yw